Tyare Corporation had the following inventory balances at the beginning and end of May:
May 1 May 30
Raw materials $ 32,500 $ 44,000
Finished Goods $ 82,000 $ 80,000
Work in Process $ 20,500 $ 17,578
During May, $65,500 in raw materials (all direct materials) were drawn from inventory and used in production. The company's predetermined overhead rate was $12 per direct labor-hour, and it paid its direct labor workers $15 per hour. A total of 440 hours of direct labor time had been expended on the jobs in the beginning Work in Process inventory account. The ending Work in Process inventory account contained $7,750 of direct materials cost. The Corporation incurred $44,100 of actual manufacturing overhead cost during the month and applied $43,800 in manufacturing overhead cost.
The amount of direct labor cost in the May 30 Work in Process inventory was:
Answer #1

Calculate direct labor cost in May 30

Ending work in process 17578
Direct material -7750
Direct labor and overhead 9828

Labor hour = 9828/27 = 364 Hour

Direct labor cost in the May 30 = 364*15 = $5460